OpenStudy (anonymous):

Rohan did a survey in four bird parks and wrote the following observations: Bird Park Survey Park Number of Yellow Birds Total Number of Birds P 33 53 Q 20 48 R 54 90 S 44 83 Based on the survey, which park had the greatest percentage of yellow birds? Park P Park Q Park R Park S

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (mathstudent55):

For each choice, divide the first number by the second number. The highest number is the answer.

OpenStudy (michele_laino):

the percentage is the ratio of number of yellow birds over number of total birds. for example in the first case, percentage is: \[\frac{ 33 }{ 53 }=0.622\] =62.2%

OpenStudy (anonymous):

ok and in the other cases?

OpenStudy (michele_laino):

please do the same with other cases!

OpenStudy (michele_laino):

in each of your cases it is a simple division

OpenStudy (michele_laino):

park Q: 20/48=0.416=41.6% please continue!

OpenStudy (michele_laino):

park R: 54/90=... park S: 44/83=...

OpenStudy (anonymous):

oh sorry, um

OpenStudy (anonymous):

what am i supposed to do?

OpenStudy (michele_laino):

park R= 60.0% park S=53.0% so now you have to pick the greates percentage

OpenStudy (michele_laino):

percentages are: park P: 62.2% park Q: 41.6% park R:60.0% park S=53.0% so?